OpenStudy (anonymous):

f(0) = -5 f(1) = 20 f(n) = -4•f(n-1) - 3•f(n-2); for n > 1

OpenStudy (anonymous):

can anyone help me figure out how to work this I got the right answer but I don't remember how I did it

OpenStudy (xapproachesinfinity):

what are you looking for here?

OpenStudy (anonymous):

just how to solve the equation again the sequence is -5, 20, -65, 200, ... Just don't remember how i figured it out

OpenStudy (xapproachesinfinity):

let me rewrite your problem \(f(0)=-5~~and ~~f(1)=20\) the recursive formula is given by \(f(n)=-4f(n-1)-3f(n-2)) ~~~for ~~n>1\)

OpenStudy (xapproachesinfinity):

so the first two term of this sequence are -5,20 and you said you are looking for how many terms?

OpenStudy (xapproachesinfinity):

is that your concern?

OpenStudy (anonymous):

I just can't seem to get the same sequence again

OpenStudy (xapproachesinfinity):

to find the next term you just use that recursive formula f(2)=-4f(1)-3f(0)=-4(20)-3(-5)=-80+15=-65

OpenStudy (anonymous):

I remember now thank you so much for your help

OpenStudy (xapproachesinfinity):

so for f(3)=-4f(2)-3f(1)=-4(-65)-3(20)=....?

OpenStudy (xapproachesinfinity):

No problem!

OpenStudy (anonymous):

it makes more sense than what I was doing

OpenStudy (xapproachesinfinity):

what were you doing?

OpenStudy (anonymous):

i was trying to plug in the numbers to figure out for 2 and using the solutions like n - (-5)

OpenStudy (xapproachesinfinity):

hmm you went off the hock lol

OpenStudy (xapproachesinfinity):

well you got it now :)

OpenStudy (anonymous):

again thank you so much

OpenStudy (xapproachesinfinity):

no problem :)